Calculate the time of free fall of an HI cloud of density 10^6
particles/m3
.
1
Expert's answer
2019-03-28T05:12:39-0400

The acceleration g felt by a test particle for a spherically symmetric distribution of mass M and radius r.

From Newton's second law, the equation of motion for a test particle at the edge of the cloud is then


"m \times g=G \times m \frac{M}{r^2} (1)"

The mass M is equal to:


"M= \frac{4 \pi}{3}\times r^3 \times ρ (2)"

where ρ is the cloud density


We put (2) in (1):


"g=\frac{4 \pi}{3}G \times r \times \rho (3)"

If it starts initially at rest, then (if acceleration is constant) it will reach the center when


"\frac{ g \times t^2}{2}=r (4)"

We put (3) in (4) and solve for t:


"t=\sqrt \frac {3}{2\times \pi \times G \times \rho} (5)"

The cloud density ρ is equal to:



"\rho=n \times mH (6)"

where n=106 particles/m3 ; mH = 1.67 × 10-27 kg


Using (6) we calculate the value of cloud density ρ: ρ= 1.67 × 10-21 kg/m3

We put the value of cloud density ρ in (5) and get:

t=2× 1015 seconds

One year contains 3× 107 seconds

In this case t=67.000.000 years


Answer: 2× 1015 seconds or 67.000.000 years